# Boosting Open Set Recognition Performance through Modulated Representation Learning

### v1

@@ Abstract
The open set recognition (OSR) problem aims to identify test samples from novel semantic classes that are not part of the training classes, a task that is crucial in many practical scenarios. However, existing OSR methods use a constant scaling factor (the temperature) to the logits before applying a loss function, which hinders the model from exploring both ends of the spectrum in representation learning – from instance-level to semantic-level features. In this paper, we address this problem by enabling temperature-modulated representation learning using our novel negative cosine scheduling scheme. Our scheduling lets the model form a coarse decision boundary at the beginning of training by focusing on fewer neighbors, and gradually prioritizes more neighbors to smooth out rough edges. This gradual task switching leads to a richer and more generalizable representation space. While other OSR methods benefit by including regularization or auxiliary negative samples, such as with mix-up, thereby adding a significant computational overhead, our scheme can be folded into any existing OSR method with no overhead. We implement the proposed scheme on top of a number of baselines, using both cross-entropy and contrastive loss functions as well as a few other OSR methods, and find that our scheme boosts both the OSR performance and the closed set performance in most cases, especially on the tougher semantic shift benchmarks [^1] The project codes are available at https://anonymous.4open.science/r/NegCosSch-4516/..
@@
Figure 1: UMAP projection of representation spaces for different temperature schedules on 10 classes of the Caltech-UCSD-Birds dataset. (a) - (c) show representations for constant temperatures ($\tau$). For lower $\tau$, the representations of open set samples are distributed and so are the representations of closed set classes, leading to a sharp decision boundary. For higher $\tau$, the representations of the known classes are more compact, making the decision boundary smoother. However, open set samples overlap with the clusters of known classes. Mid value of $\tau$ achieves a trade-off but does not gain the benefits of both ends. (e) - (g) show representations of our temperature scheduling scheme NegCosSch as the training progresses. A lower $\tau$ at the start leads to a coarse decision and the model gradually makes the classes more compact and the open set representations are pushed away. Finally, (d) show the representation space for a previous temperature scheduling scheme CosSch, which is better than constant temperatures but not as compact as our NegCosSch. Experimental details appear in the Appendix.
1 Introduction
@@
Early research attempts model the unknown classes as a long-tailed distribution using extreme value theory (Vignotto and Engelke (2018)) to modify the softmax function (Bendale and Boult (2016)), while other methods generate unknown samples either with mix-up or using generative models to represent novel classes (Zhou et al. (2021); Ge et al. (2017); Neal et al. (2018); Chen et al. (2021)). The synthetic examples may not generalize well to a variety of unknown classes. Later methods add regularization to explicitly bound the risks for novel classes (Zhou et al. (2021); Chen et al. (2021, 2020a)). In essence, these methods create more empty regions in the representation space hoping that unknown representations lie in those regions. The tougher unknown samples lie very close to the known classes in representation space, which reduce the effectiveness of such methods.
@@
Vaze et al. (2022) establish new OSR baselines by training models with better design choices and argue that well-trained closed set classifiers achieve better OSR performance, where the unknown samples exhibit lower max logit scores. This essentially has motivated the next generation of OSR methods to learn even better representations for improving OSR performance. The recent methods essentially aim for better representation learning with a better loss function, such as the contrastive loss (Khosla et al. (2020); Chen et al. (2020b)) with sample mix-up (Verma et al. (2018); Zhang et al. (2017)) and by adding different regularization schemes (Zhou et al. (2024); Bahavan et al. (2025); Li et al. (2025); Wang et al. (2025)).
@@
Moreover, the regular OSR benchmarks commonly used are small in scale. In this regard, semantic shift benchmarks (SSBs) are proposed by Vaze et al. (2022) on three fine-grained datasets, having more classes with varying levels of OSR difficulty. Therefore, the OSR methods that demonstrate improvement on smaller datasets but involve either data generation (Chen et al. (2021); Ge et al. (2017)), mix-up (Zhou et al. (2021); Xu et al. (2023)), or training an additional model with separate objective, such as VAEs that include reconstruction based objective (Oza and Patel (2019); Yoshihashi et al. (2019); Zhou et al. (2024)), are computationally demanding and often require higher memory. These methods are unsuitable for SSBs as the base model for SSBs requires a significant compute for good closed set performance. Most of the latest research does not use these benchmarks. Therefore, better representation learning schemes are required, which add little to no computational overhead.
@@
To achieve this, we need to explore the inner mechanisms of the losses that are the basis of most OSR methods, such as the cross-entropy (CE) loss and the contrastive loss. These loss functions compute probabilities by applying a temperature scaling to the logits- the model’s raw outputs- where the temperature coefficient adjusts the sharpness of the resulting probability distributions. Temperature is the key parameter to control the learned features for both losses. Works by Zhang et al. (2022); Kukleva et al. (2023); Zhang et al. (2021) demonstrate that learning instance-specific representations are encouraged by a lower temperature value, while semantic representations are learned through a higher value. However, a fixed temperature throughout the training prevents the model from exploring both ends of this learning spectrum. In this regard, Kukleva et al. (2023) demonstrate the benefits of learning both instance-level and semantic features in self-supervised training for long-tailed datasets in closed set scenarios using a cosine temperature scheduling.
@@
In OSR, learning a representation space that provides both instance-specific and semantic features is also crucial to achieve better open set and closed set performance. In this research, we analyze the representation space for different temperature values on both losses. Based on our analysis, we formulate a novel temperature scheduling scheme for temperature modulated representation learning. We find temperature modulation with the proposed scheduling is beneficial to create more compact clusters for representing closed set classes, while keeping more open set examples distant from these clusters, resulting in overall better representations.
@@
  • We propose a novel temperature scheduling (TS) scheme called the negative cosine scheduling (NegCosSch) for temperature modulated representation learning. The scheme can be seamlessly integrated into any existing OSR algorithm, with both the CE and the losses based on contrastive learning, without any computational overhead. We show performance improvements on both the regular OSR benchmarks and the SSBs.
  • We further integrate our method with a few existing methods for OSR and show a significant boost in their performance.
  • Our scheduling demonstrates strong performance improvements for the tougher SSBs over the baselines on both the closed set and the open set problems. We show that our scheme brings out stronger improvements with an increased number of training classes when the task becomes more difficult for the baseline model.
The rest of the paper is organized as follows. In Section 2, we discuss the relevant background with different losses. In Section 3, we describe the proposed TS scheme. In Section 4, we discuss our results followed by the related works in Section 5 and the concluding remarks in Section 6.
2 Background
During training, the model is decomposed into two components: The first component is an encoder function $f(\cdot)$ which maps the input $x$ to a representation $z=f(x)$. The second component $h(\cdot)$ maps the representations to task specific outputs, which is either a linear classification layer if we train with the CE loss or a projection layer if we use the contrastive training. The final outputs, also called the logits, $l=h(z)=h(f(z))$ are then given as model predictions to the loss functions. We assume for a specific problem, a model is trained for a predefined number of epochs $E$. We further discuss the CE and the supervised contrastive (SupCon) losses and the effects of the temperature parameter, which provide a basis for our proposed scheme.
@@ 2.1 Cross-entropy Loss
For a batch of training data =$\mathcal{B}=\{(x_{k},y_{k})\}_{k=1}^{B}$, the CE loss is calculated as
@@
[…]
@@ 2.2 Supervised Contrastive Loss
For a given batch $\mathcal{B}$, the SupCon training utilizes a multi-viewed batch by taking two augmented samples of the same original sample. The multi-viewed batch =$\mathcal{B}^{\prime}=\{(\tilde{x_{i}},\tilde{y}_{i})\}_{i=1}^{2B}$, where $\tilde{x}_{2k}$ and $\tilde{x}_{2k-1}$ are two random augmentations of $x_{k}~{}~{}(k=1,...,B)$, and $\tilde{y}_{2k-1}=\tilde{y}_{2k}=y_{k}$. Below we refer to $i$ as the anchor index from $I=\{1,...,2B\}$. The SupCon loss (Khosla et al. (2020)) is defined by:
@@
Here, $\text{sim}(l_{i},l_{j})$ is the cosine similarity between $l_{i}$ and $l_{j}$. $P(i)=\{p\neq i:\tilde{y}_{i}=\tilde{y}_{p}\}$ is the set of indices of all samples having the same label as $\tilde{y}_{i}$ in $\mathcal{B}^{\prime}$ distinct from $i$. Contrastive loss, by construction, gains its strength by pushing away the representations of the negative samples (samples of other classes) and by producing compact clusters of representations for the (positive) samples of the same class.
2.3 Effect of Temperature on the Losses
The SupCon loss applies hard negative mining from penalizing the harder negative samples more through the exponential function (Khosla et al. (2020)). The measure of hardness of a sample with respect to an anchor sample is determined by the similarity function. Therefore as a scaling factor, the temperature plays a critical role in controlling the trade-off between uniformity and semantic structure in the representation space as shown in Wang and Liu (2021); Kukleva et al. (2023) for the self-supervised loss (Chen et al. (2020b)). This effect mostly translates to the supervised case except for the fact that the definition of positive and negative samples are now different.
For smaller values of temperature $\tau$, as the differences in similarity get amplified, the nearest negative samples receive the highest gradient (Wang and Liu (2021)). The model aggressively pushes the negative samples away, leading to features that are appropriate for instance-level discrimination and distributing the embeddings over the representation space. However, the positive samples do not cluster tightly because, like the negative samples, fewer positive neighbors get priority in the loss function (Figure 1(a)). The resulting decision boundary is sharper. The open set samples do not get closer to the known classes due to the heavy penalty of having slight dissimilarity.
With larger $\tau$, the differences in similarity diminish and the repulsive force gets distributed to more negative neighbors. The model can decrease the loss by learning the group-wise features rather than the instance discriminating features to push away easy negatives, inducing semantic structures. Due to compact clusters of within-class representations, the resulting decision boundary is smoother. However, the between-class distances also reduce as the model is less aggressive in removing the negatives. In this case a lot of open set examples get close to the known classes (Figure 1(c)).
@@ 3 Proposed Method
In this section, we formally introduce our problem, describe our proposed TS scheme and how the scheduling leads to learning representations useful for OSR.
@@ 3.1 Problem Definition
We are given a labeled training dataset $\mathcal{D}_{tr}=\{(x_{i},y_{i})\}\subset\mathcal{X}\times\mathcal{Y}$, where $x_{i}$ is the training sample with label $y_{i}$. $\mathcal{X}$ is the input space and the labels of $\mathcal{D}_{tr}$ from a closed set label space $\mathcal{Y}$, i.e. $y_{i}\in\mathcal{Y},\forall i$. The total number of classes in the closed set is $C=|\mathcal{Y}|$. The test dataset $\mathcal{D}_{test}\subset\mathcal{X}\times\mathcal{Y}\cup\mathcal{O}$ consists of samples whose label space $\mathcal{Y}\cup\mathcal{O}$ is different than $\mathcal{Y}$, and $\mathcal{O}\cap\mathcal{Y}=\emptyset$. $\mathcal{O}$ is the set of unknown classes termed as the open set. The objective of OSR is to classify a test sample among the closed set classes or to flag it as belonging to an unknown class. We assume that information about the nature of unknown classes or any auxiliary samples are unavailable during training.
Figure 2: Different temperature scheduling schemes.
3.2 Proposed Temperature Scheduling
For a temperature dependent loss function, $\tau$ is usually kept constant throughout the training process. We denote constant TS by Const, defined as
$\text{Const}(e;\tau)=\tau,\forall e$ (3)
@@ 3.2 Proposed Temperature Scheduling
For OSR, if the test representation from a novel class is too class-specific, the model easily finds its similarity to one of the known classes. On the other hand, if the feature is too instance-specific, the model is under-confident in assigning any sample to a known class. Therefore, a model needs to capture both features – good class-specific representations having room for instance level discriminating power within the class. A specific temperature throughout the training restricts the opportunity to explore this trade-off that may lead to learning better representations. We propose to gradually switch between the two objectives using a generalized cosine scheduling, which we describe next.
3.2.1 Generalized Cosine Scheduling

### v2

@@ Abstract
The open set recognition (OSR) problem aims to identify test samples from novel semantic classes that are not part of the training classes, a task that is crucial in many practical scenarios. However, the existing OSR methods use a constant scaling factor (the temperature) to the logits before applying a loss function, which hinders the model from exploring both ends of the spectrum in representation learning – from instance-level to semantic-level features. In this paper, we address this problem by enabling temperature-modulated representation learning using a set of proposed temperature schedules, including our novel negative cosine schedule. Our temperature schedules allow the model to form a coarse decision boundary at the beginning of training by focusing on fewer neighbors, and gradually prioritizes more neighbors to smooth out the rough edges. This gradual task switching leads to a richer and more generalizable representation space. While other OSR methods benefit by including regularization or auxiliary negative samples, such as with mix-up, thereby adding a significant computational overhead, our schedules can be folded into any existing OSR loss function with no overhead. We implement the novel schedule on top of a number of baselines, using cross-entropy, contrastive and the ARPL loss functions and find that it boosts both the OSR and the closed set performance in most cases, especially on the tougher semantic shift benchmarks. Project codes will be available.
1 Introduction
@@
Figure 1: UMAP projection of representation spaces for different temperature schedules on 10 classes of the Caltech-UCSD-Birds dataset. (a)- (c) show representations for constant temperatures ($\tau$). For lower $\tau$, the representations of unknown classes are distributed and so are the representations of known classes, leading to a sharp decision boundary. For higher $\tau$, the representations of known classes are more compact, making the decision boundary smoother. However, unknown samples overlap with the clusters of known classes. Mid value of $\tau$ achieves a trade-off but does not gain the benefits of both ends. (e)- (g) show representations of our temperature schedule NegCosSch as the training progresses. A lower $\tau$ at the start leads to a coarse decision and the model gradually makes the classes more compact and the unknown representations are pushed away. Finally, (d) show the representation space for a previous schedule CosSch, which is better than fixed temperatures but not as compact as our NegCosSch. Experimental details appear in the Appendix.
@@
Most of the early research attempts either model the unknown classes as long-tailed distributions (Vignotto & Engelke, 2018; Bendale & Boult, 2016), generate synthetic samples using generative models (Ge et al., 2017; Neal et al., 2018; Chen et al., 2021; Moon et al., 2022) or with mix-up (Chen et al., 2021; Xu et al., 2023; Li et al., 2024; Zhou et al., 2021) to represent novel classes, or train a secondary model with a separate objective, such as VAEs that include reconstruction based objective (Oza & Patel (2019); Yoshihashi et al. (2019); Zhou et al. (2024a)). The synthetic examples may not generalize well to a variety of unknown classes, whereas training generative or secondary models or with mix-up are computationally demanding and often require higher memory. Later methods add regularization (Zhou et al., 2021; Chen et al., 2021; 2020a) to explicitly bound the open space risks. In essence, these methods create more empty regions in the representation space hoping that unknown representations lie in those regions. Forcing the creation of empty spaces does not result in an improved OSR as the tougher unknown samples lie close to known classes in representation space, incurring significant similarity between them and reducing the effectiveness of such methods.
@@
Vaze et al. (2022) establish new OSR baselines by training models with optimal design choices and argue that a well-trained closed set classifier achieves an improved OSR performance, where the unknown samples exhibit lower max-logit scores. This essentially has motivated the next generation of OSR methods to learn even better representations for improving performance through a better loss function, such as the contrastive loss (Khosla et al. (2020); Chen et al. (2020b)) with sample mix-up (Xu et al. (2023); Verma et al. (2018); Zhang et al. (2017)) and by adding different regularization schemes (Zhou et al. (2024a); Bahavan et al. (2025); Li et al. (2025); Wang et al. (2025)).
@@
Moreover, the regular OSR benchmarks commonly used are small in scale. In this regard, semantic shift benchmarks (SSBs) are proposed by Vaze et al. (2022) on fine-grained datasets, having more classes with varying levels of OSR difficulty. Therefore, the methods that demonstrate improvement on smaller datasets but involve either data generation, mix-up, or training secondary models are unsuitable for the larger benchmarks as training a well-performing base model on them requires a significant compute and memory. Most of the latest research does not use these benchmarks. This necessitates the development of an advanced representation learning scheme that impose minimal computational overhead.
@@
To achieve this, we need to explore the inner mechanisms of the losses that are the basis of most OSR methods, such as the cross-entropy (CE) and the contrastive loss. These loss functions compute probabilities by applying a temperature scaling to the logits- the model’s raw outputs- where the temperature coefficient adjusts the sharpness of resulting probability distributions. It is the key parameter to control the learned features for both losses. Prior works (Wang & Liu, 2021; Zhang et al., 2022; Kukleva et al., 2023; Zhang et al., 2021) demonstrate that a lower temperature encourages instance-specific representations while a higher value encourages semantic ones. However, a fixed temperature throughout the training prevents the model from exploring both ends of this learning spectrum. In this regard, Kukleva et al. (2023) study the benefits of learning both instance-level and semantic features primarily in the closed-set scenarios using self-supervised contrastive learning for long-tailed datasets using a cosine temperature schedule (TS), but the impact of temperature scaling or a TS remains largely unexamined for novel classes.
@@
In OSR, learning a representation space that provides both instance-specific and semantic features is also crucial to achieve improved open set and closed set performance. In this research, we analyze the representation space for different temperature scaling factors on both losses in an open set scenario. Based on the analysis, we propose novel temperature schedules for temperature modulated representation learning. We find temperature modulation with the proposed schedules is beneficial to create more compact clusters for representing the closed set classes, while keeping open set examples more distant from these clusters, resulting in overall improved representations.
@@
  • We analyze the effects of temperature scaling in an open set scenario, using a number of TSs, including our novel negative cosine schedule (NegCosSch), to explore temperature modulated representation learning. We find that the proposed schedules, even simple linear schedules, demonstrate better open and closed set performance compared to the usual constant temperature baselines and possible other schedules.
  • Our schedules can be seamlessly integrated into any existing OSR loss, such as the CE, the losses based on contrastive learning and the ARPL loss by Chen et al. (2021), without any computational overhead. We show significant performance improvements on the TinyImageNet benchmark and the SSBs.
  • Our strategy demonstrates strong performance improvements for the tougher SSBs over the baselines for both the closed set and the open set problems. We show that our scheme achieves stronger improvements with an increased number of training classes when the task becomes more difficult for the baseline model.
The rest of the paper is organized as follows. In Section 2, we discuss the relevant background on different losses and in Section, 3, we discuss the effect of temperature scaling on known and unknown classes. In Section 4, we describe the proposed scheme. In Section 5, we discuss our results followed by related works in Section 6 and present the concluding remarks in Section 7.
2 Background on Losses
During training, the model is decomposed into two components: The first component is an encoder function $f(\cdot)$ which maps the input $x$ to a representation $z=f(x)$. The second component $h(\cdot)$ maps the representations to task specific outputs, which is either a linear classification layer if we train with the CE loss or a projection layer if we use the contrastive training. The final outputs, also called the logits, $l=h(z)=h(f(z))$ are then given as model predictions to the loss functions. We assume for a specific problem, a model is trained for a predefined number of epochs $E$. We further discuss the CE and the supervised contrastive (SupCon) losses and the effects of the temperature parameter, which provide a basis for our proposed scheme. The ARPL loss is discussed in Appendix E.
@@ 2.1 Cross-entropy Loss
For a batch of training data $\mathcal{B}=\{(x_{k},y_{k})\}_{k=1}^{B}$, the CE loss is calculated as
@@
where $p_{k}=\text{softmax}(l_{k}/\tau)$, $\text{one\_hot}(y_{k})$ is the one hot encoded vector of $y_{k}$ and $(\cdot)$ is the dot product. The parameter $\tau>0$ is called the temperature.
@@ 2.2 Supervised Contrastive Loss
For a given batch $\mathcal{B}$, the SupCon training utilizes a multi-viewed batch by taking two augmented samples of the same original sample. The multi-viewed batch $\mathcal{B}^{\prime}=\{(\tilde{x_{i}},\tilde{y}_{i})\}_{i=1}^{2B}$, where $\tilde{x}_{2k}$ and $\tilde{x}_{2k-1}$ are two random augmentations of $x_{k}~~(1\leq k\leq B)$, and $\tilde{y}_{2k-1}=\tilde{y}_{2k}=y_{k}$. Below we refer to $i$ as the anchor index from $I=\{1,...,2B\}$. The SupCon loss (Khosla et al. (2020)) is defined by:
@@
Here, $\text{sim}(l_{i},l_{j})$ is the cosine similarity between $l_{i}$ and $l_{j}$. $P(i)=\{p\neq i:\tilde{y}_{i}=\tilde{y}_{p}\}$ is the set of indices in $\mathcal{B}^{\prime}$ having the same label as $\tilde{y}_{i}$ and distinct from $i$. Contrastive loss, by construction, gains its strength by pushing away the representations of the negative samples (samples of other classes) and by producing compact clusters of representations for the (positive) samples of the same class.
3 Effect of Temperature on Known and Unknown Samples
The SupCon loss applies hard negative mining from penalizing the harder negative samples more through the exponential function (Khosla et al. (2020)). The measure of hardness of a sample with respect to an anchor is determined by the scaled similarity. Therefore as a scaling factor, the temperature plays a critical role in controlling the trade-off between uniformity and semantic structure in the representation space as shown in Wang & Liu (2021); Kukleva et al. (2023) for the self-supervised loss (Chen et al. (2020b)). This effect mostly translates to the supervised case except for the fact that the definition of positive and negative samples are now different.
For any given anchor index $i$, the gradient of $L_{SupCon}$ with respect to a negative logit $l_{j}$ can be computed as shown in the following equation.
$\frac{\partial L_{SupCon}}{\partial l_{j}}=\frac{\partial L_{SupCon}}{\partial\text{sim}({l_{i},l_{j}})}\times\frac{\partial\text{sim}({l_{i},l_{j}})}{\partial l_{j}}=\frac{1}{\tau}[\text{softmax}_{a\in I\setminus\{i\}}(\text{sim}(l_{i},l_{a})/\tau)]_{j}\times\frac{\partial\text{sim}({l_{i},l_{j}})}{\partial l_{j}}$ (3)
For smaller values of temperature $\tau$, as the differences in scaled similarity get amplified, the nearest negative samples receive the highest gradient (Wang & Liu (2021)) and the model minimizes similarity to them with respect to anchor $i$. The model aggressively pushes the nearest negative samples away, leading to features that are appropriate for instance-level discrimination and distributing the embeddings over the representation space. However, the positive samples do not cluster tightly because, like the negative samples, fewer positive neighbors get priority in the loss function (Figure 1(a)). The resulting decision boundary is sharper. The open set representations do not get closer to the known classes due to the heavy penalty of having slight dissimilarity.
With larger $\tau$, the differences in scaled similarity diminish and the repulsive force gets distributed to more negative neighbors. The model can decrease the loss by learning the group-wise features rather than the instance discriminating features to push away easy negatives, inducing semantic structures. Due to compact clusters of within-class representations, the resulting decision boundary is smoother. However, the between-class distances also reduce as the model is less aggressive in removing the negatives. In this case a lot of open set examples get close to the known classes (Figure 1(c)).
@@ 4 Proposed Method
In this section, we formally introduce our problem, describe our proposed temperature modulation and explain how our schedules lead to learning representations useful for OSR.
@@ 4.1 Problem Definition
We are given a labeled training dataset $\mathcal{D}_{tr}=\{(x_{i},y_{i})\}\subset\mathcal{X}\times\mathcal{Y}$, where $x_{i}$ is the training sample with label $y_{i}$. $\mathcal{X}$ is the input space and the labels of $\mathcal{D}_{tr}$ come from a closed set label space $\mathcal{Y}$, i.e. $y_{i}\in\mathcal{Y},\forall i$. The total number of classes in the closed set is $C=|\mathcal{Y}|$. The test dataset $\mathcal{D}_{test}\subset\mathcal{X}\times\mathcal{Y}\cup\mathcal{O}$ consists of samples whose label space $\mathcal{Y}\cup\mathcal{O}$ is different than $\mathcal{Y}$, and $\mathcal{O}\cap\mathcal{Y}=\emptyset$. $\mathcal{O}$ is the set of unknown classes defined as the open set. The objective of OSR is to classify a test sample among the closed set classes or to flag it as belonging to an unknown class. We assume that information about the nature of unknown classes or any auxiliary samples are unavailable during training.
Figure 2: Different temperature schedules.
4.2 Rationale for Temperature Scheduling
For a temperature dependent loss function, $\tau$ is usually kept constant throughout the training process. We denote a constant TS by $\mathcal{T}_{\text{Const}}$, defined as
$\mathcal{T}_{\text{Const}}(e;\tau)=\tau,\forall e$ (4)
@@ 4.2 Rationale for Temperature Scheduling
For OSR, if the test representation from a novel class is too class-specific, the model easily finds its similarity to one of the known classes. On the other hand, if the feature is too instance-specific, the model is under-confident in assigning any sample to a known class. Moreover, Familiarity Hypothesis by Dietterich & Guyer (2022) states that most existing OSR methods flag semantic novelty from the absence of learned semantic features and it recommends to extract features for interesting content beyond the semantic features for detecting novelty. Therefore, a model needs to capture both features – good class-specific representations while having room for instance-level discriminating power within the class. If for the group-wise features, there is similarity between a novel sample and a known class, the instance-specific features should maintain the separation between them. A specific temperature throughout the training restricts the opportunity to explore this trade-off that may lead to learning better representations. Utilizing the effects observed in Section 3, we propose to gradually switch between the two objectives using a generalized cosine schedule, which we describe next.
4.3 Proposed Temperature Schedules